OpenStudy (anonymous):

Please help!! How do you solve 9y^0, I have never seen the equation with a 0

OpenStudy (anonymous):

Is it (9y)^0 or 9y^0

OpenStudy (anonymous):

9y^0

OpenStudy (anonymous):

9y^0 = 9

OpenStudy (anonymous):

Because y^0 is 1.

OpenStudy (anonymous):

if y =1 then wouldn't it stay as 9y?

OpenStudy (lalaly):

y^0=1 not y=1

OpenStudy (anonymous):

yes, y^0 = 1.

OpenStudy (lilg132):

9y^0 = 9 ??

OpenStudy (anonymous):

any number to the power of zero is equal to 1

OpenStudy (lalaly):

yeah^

OpenStudy (anonymous):